I am taking college Algebra this year and I forgot how to solve this problem may you please help me.
The problem is this: 7/8+i.
In the instructions it says that I have to write it in the form A+Bi

Answer by Alan3354(69443)

7/8+i
-------
If you mean 7/8 + i, it's done.
If you mean 7/(8+i)
Multiply NUM and DEN by the conjugate of the DEN, 8-i
----
= 7(8-i)/(64-i^2)
= (56 - 7i)/65
= 56/65 - 7i/65